if a squared minus b squared is equal to 8 and a minus b is equal to 2 ,find the value of (a+b)squared
Answer by rothauserc(4718)

(a^2 - b^2) = 8
factor polynomial
(a+b) * (a - b) = 8
we know (a - b) = 2, therefore
(a + b) * 2 = 8
divide both sides of = by 2
(a + b) = 4
and
(a + b)^2 = 16